# Towards Global Localization using Multi-Modal Object-Instance Re-Identification

### v1

@@ V Results
CLIP-ReiD (CNN) [li2023clipreid] 54.1
TransReID [he2021transreid] 55.7
LoGoViT [10096126] 60.2
CLIP-ReID (ViT) [li2023clipreid] 61.2
NFormer [wang2022nformer] 63.2
PADE [wang2024parallel] 66.0
@@
Table II: ReID mAP Metrics. We benchmark DATOR against several ReID methods on our DATOR-ReID dataset. We show competitive results with only one modality and significantly outperform with both. Best result is in bold and second-best is underlined.
We benchmark DATOR against some popular vehicle-ReID and person-ReID baselines, the results are presented in table II. We train all the RGB-only baselines on the RGB images in the DATOR-ReID dataset. To evaluate the effectiveness of each individual pathway in DATOR, we also show RGB-only inference results, in which the depth features are zeroed (similar to modality dropout) and depth-only inference results, wherein the RGB features are zeroed. Note that DATOR achieves competitive results even when using only one of the modalities, while at the same time achieving much better performance (last row in table II) when both the modalities are used. This shows that there is effective exchange of information between both the pathways in the network.
Further, we show a qualitative result in figure 5 where DATOR re-identifies the correct instance in a low-illumination setting while the existing best performing method, PADE [wang2024parallel], retrieves a similar but incorrect instance.
@@
Tables I(c), I(c) and I(c) summarize the effectiveness of each model in our localization architecture. We include both mean and median errors to illustrate the effect of high error misalignments as compared to successful alignments, that are near perfect predictions of pose. The large jump in success rate from other baselines to DATOR points to more successful instance ReID leading to more accurate registration. Our results demonstrate the challenging nature of our datasets and the potential improvements left to future works. A common trend observed across all models is the negative effect of large, flat, texture-less walls that interfere with ICP matching. For example, unlike the other TUM sequences, fr1_room dataset from TUM has large, textureless, flat objects that result in low success rates for all benchmarked methods.
In comparison, results shown in [matsuzaki2024clip] are only for two TUM sequences (fr2_desk, fr3_lo__household.) achieving close to only $80$% while we achieve success rates of $88.5$% and $100$% respectively. This is despite of us employing stricter margins and also considering rotation error (See Fig. 4 in [matsuzaki2024clip]).
Figure 5: Qualitative Analysis of DATOR. Given a query in a low-illumination scene, DATOR reidentifies the robot instance successfully, while PADE, identifies it as a different robot that is missing an overhead attachment. This gain can be attributed to DATOR’s use of depth information.
VI Conclusion

### v2

@@ 4. Experimental Setup
CLIP-ReiD (CNN) (li2023clipreid, ) 54.1
TransReID (he2021transreid, ) 55.7
LoGoViT (10096126, ) 60.2
CLIP-ReID (ViT) (li2023clipreid, ) 61.2
NFormer (wang2022nformer, ) 63.2
PADE (wang2024parallel, ) 66.0
@@
Table 3. ReID mAP Metrics. We benchmark DATOR against several ReID methods trained our DATOR-ReID dataset. We demonstrate the effectiveness of using both modalities for object ReID. Best result is in bold and second-best is underlined.
Object ReID: DATOR was trained for 240 epochs, using SGD optimizer and cosine LR scheduler, with an initial learning rate of 0.008. A batch size of 64 was used for all the experiments.
Localization: We evaluate our localization pipeline on DATOR-lab, 6 DATOR-synth sequences, as well as 5 TUM RGB-D sequences, (namely fr{_desk,_desk2,_room}, freiburg2_desk and
freiburg3 _long_office_household).
For DATOR-synth, we create an object memory for each sequence by sampling RGB-D pairs every 30 frames. Additional unseen RGB-D pairs are sampled every 30 frames with a 15-frame offset. The only information available during localization is the generated object memory. Each frame is individually localized using the sequence’s object memory. Sampling strategies for other datasets are detailed in our released implementation.
We compare each RGB-D pair’s estimated pose with its corresponding ground truth pose, measuring 3D translation error (TE) in meters and rotation error (RE) in radians. An RGB-D pair is said to be correctly localized, if the TE and RE are below certain thresholds (see caption under table 2). We report the success rates for various baselines along with our method for comparison.
The implementations and annotation tools used for (zhang2023instaloc, ) and (matsuzaki2024clip, ) are not available. We release a re-implementation of (matsuzaki2024clip, ) but observe suboptimal results using this approach (elaborated in section 5).
5. Results
Object ReID Results
We benchmark DATOR against some popular vehicle-ReID and person-ReID baselines, the results are presented in table 3. We train all the RGB-only baselines on the RGB images in the DATOR-ReID dataset. To evaluate the effectiveness of each individual pathway in DATOR, we also show RGB-only inference results, in which the depth features are zeroed (similar to modality dropout) and depth-only inference results, wherein the RGB features are zeroed. Note that DATOR achieves competitive results even when using only one of the modalities, while at the same time achieving much better performance (last row in table 3) when both the modalities are used. This shows that there is effective exchange of information between both the pathways in the network.
Further, we show a qualitative result in Fig. 4 where DATOR re-identifies the correct instance in a low-illumination setting while the existing best performing method, PADE (wang2024parallel, ), retrieves a similar but incorrect instance.
@@ 5. Results
Table 2 summarizes the effectiveness of each model in our localization architecture. We include both mean and median errors to illustrate the effect of high error misalignments as compared to successful alignments, that are near perfect predictions of pose. The large jump in success rate from other baselines to DATOR points to more successful instance ReID leading to more accurate registration. Our results demonstrate the challenging nature of our datasets and the potential improvements left to future works. A common trend observed across all models is the negative effect of large, flat, texture-less walls that interfere with ICP matching. For example, unlike the other TUM sequences, fr1_room dataset from TUM has large, textureless, flat objects that result in low success rates for all benchmarked methods.
In comparison, results shown in (matsuzaki2024clip, ) are only for two TUM sequences (fr2_desk, fr3_lo__household.) achieving close to only $80$% while we achieve success rates of $88.5$% and $100$% respectively. This is despite of us employing stricter margins and also considering rotation error (See Fig. 4 in (matsuzaki2024clip, )).
Figure 4. Qualitative Analysis of DATOR. Given a query in a low-illumination scene, DATOR reidentifies the robot instance successfully, while PADE, identifies it as a different robot that is missing an overhead attachment. This gain can be attributed to DATOR’s use of depth information.
6. Conclusion